# Maximal Cohen-Macaulay DG-Complexes

### v1

@@ 3.3. DG-Modules of Maximal Depth
(3.1) ${R}$${\Sigma^{n}M}$${k}$${k\otimes_{R}\Sigma^{n}M}$$\scriptstyle{\cdot z}$
where the dashed line is a left inverse of the arrow above it. To see that the dashed line exists, consider the map $k\to k\otimes_{R}\Sigma^{n}M$. There is a nonzero splitting ${k}$${\operatorname{H}(k\otimes_{R}\Sigma^{n}M)}$ in $\mathsf{D}(k)$ since the image of $z$ in $k\otimes_{R}M$ is not a boundary. Since $\operatorname{H}(k\otimes_{R}\Sigma^{n}M)\cong k\otimes_{R}\Sigma^{n}M$ in $\mathsf{D}(k)$, we have a splitting ${k}$${k\otimes_{R}\Sigma^{n}M}$ which lifts to $\mathsf{D}(R)$ through restriction of scalars. By applying the $F\otimes^{\mathbf{L}}_{R}-$ functor and taking the zeroth cohomology to Diagram 3.1, we obtain the commutative diagram
(3.2) ${\operatorname{H}^{0}(F)}$${\operatorname{H}^{0}(F\otimes_{R}\Sigma^{n}M)}$${\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)}$${\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}(k\otimes_{R}\Sigma^{n}M))}$
@@
$\geq\dim\operatorname{H}^{0}(R)_{\mathfrak{p}}-s$
Note that $0pt_{R_{\mathfrak{p}}}(F\otimes_{R}M)_{\mathfrak{p}}=s$ by Lemma 3.1 since $\mathfrak{p}R_{\mathfrak{p}}$ is an associated prime of $\operatorname{H}^{s}(F\otimes_{R}M)_{\mathfrak{p}}$ , and that Equation 3.3 is from Lemma 3.7.
@@
This implies a contradiction, and so $\mathfrak{p}=\mathfrak{m}$. As a result, Equation 3.3 becomes
@@
Let $s=n$. Since $\operatorname{H}^{0}(F)$ is finitely generated, every minimal generator in $\operatorname{H}^{0}(F)$ gives rise to a minimal generator in $\operatorname{H}^{0}(F)\otimes_{\operatorname{H}^{0}(R)}k$ by Nakayama’s Lemma. By [3, Lemma 3.2], $\operatorname{H}^{0}(F)\otimes_{\operatorname{H}^{0}(R)}k\cong\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)$ as $\sup(F)=0$. The minimal generator in $\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)$ gives rise to a nonzero element in $\operatorname{H}^{n}(F\otimes^{\mathbf{L}}_{R}(k\otimes_{R}M)$ by Diagram 3.2 in Theorem 3.5. The commutativity of Diagram 3.2 implies that there is a nonzero element in $\operatorname{H}^{n}(M\otimes_{R}F)$ that is annihilated by $I$, and so $\Gamma_{I}\operatorname{H}^{n}(M\otimes_{R}F)\neq 0$. Since $\inf(M\otimes_{R}F)=n$, this implies that $0pt_{R}(I,M\otimes_{R}F)=n$.
@@ 4.1. Definition
(4.1) $\operatorname{Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{n}((k\otimes^{\mathbf{L}}_{R}M)^{{\dagger}{\dagger}}),\operatorname{H}^{0}(E))\to\operatorname{Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{n}(M^{{\dagger}{\dagger}}),\operatorname{H}^{0}(E))$
which is nonzero since $\operatorname{H}^{0}(E)$ is a faithfully injective $\operatorname{H}^{0}(R)$-module. By [16, Thm. 4.10], Equation 4.1 can be rewritten as
(4.2) $\operatorname{H}^{-n}(\operatorname{\mathbf{R}Hom}_{R}((k\otimes_{R}M)^{{\dagger}{\dagger}},E))\to\operatorname{H}^{-n}(\operatorname{\mathbf{R}Hom}_{R}((M)^{{\dagger}{\dagger}},E))$
@@
The third isomorphism holds since $E\cong\Sigma^{d}\operatorname{\mathbf{R}\Gamma}_{\mathfrak{m}}(D)$ by [16, Prop. 7.25], and the fifth isomorphism holds by [16, Prop. 7.24] and Greenless-May duality. The sixth isomorphism holds since the DG-module $k$ is $\mathfrak{m}$-torsion. Equation 4.2 now becomes
(4.3) $\operatorname{H}^{-n}(\Sigma^{d}\operatorname{\mathbf{R}Hom}_{R}(k,M^{{\dagger}}))\to\operatorname{H}^{-n}(\Sigma^{d}(\operatorname{\mathbf{R}\Gamma}_{\mathfrak{m}}(M^{{\dagger}})))$
@@ 5. The Canonical Element Theorem for DG-Rings
(5.1) $\phi:\operatorname{\mathbf{R}Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{0}(R)\otimes^{\mathbf{L}}_{R}k,\operatorname{H}^{0}(D))\to\operatorname{\mathbf{R}Hom}_{\operatorname{H}^{0}(R)}(k,\operatorname{H}^{0}(D))$
@@
(5.2) $\xi_{\operatorname{H}^{0}(D)}^{d}:\operatorname{Ext}_{R}^{d}(k,\operatorname{H}^{0}(D))\to\operatorname{H}_{\mathfrak{m}}^{d}(\operatorname{H}^{0}(D))$

### v2

@@ Abstract.
Part of this article was written while the author was supported by the “RTG: Commutative Algebra at Nebraska” grant (NSF grant 2342256).
@@ 3.3. DG-Modules of Maximal Depth
(3.3.1) ${R}$${\Sigma^{n}M}$${k}$${k\otimes_{R}\Sigma^{n}M}$$\scriptstyle{\cdot z}$
where the dashed line is a left inverse of the arrow above it. To see that the dashed line exists, consider the map $k\to k\otimes_{R}\Sigma^{n}M$. There is a nonzero splitting ${k}$${\operatorname{H}(k\otimes_{R}\Sigma^{n}M)}$ in $\mathsf{D}(k)$ since the image of $z$ in $k\otimes_{R}M$ is not a boundary. Since $\operatorname{H}(k\otimes_{R}\Sigma^{n}M)\cong k\otimes_{R}\Sigma^{n}M$ in $\mathsf{D}(k)$, we have a splitting ${k}$${k\otimes_{R}\Sigma^{n}M}$ which lifts to $\mathsf{D}(R)$ through restriction of scalars. By applying the $F\otimes^{\mathbf{L}}_{R}-$ functor and taking the zeroth cohomology to Diagram 3.3.1, we obtain the commutative diagram
(3.3.2) ${\operatorname{H}^{0}(F)}$${\operatorname{H}^{0}(F\otimes_{R}\Sigma^{n}M)}$${\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)}$${\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}(k\otimes_{R}\Sigma^{n}M))}$
@@
(3.3.3) $\geq\dim\operatorname{H}^{0}(R)_{\mathfrak{p}}-s$
Note that $0pt_{R_{\mathfrak{p}}}(F\otimes_{R}M)_{\mathfrak{p}}=s$ by Lemma 3.1 since $\mathfrak{p}R_{\mathfrak{p}}$ is an associated prime of $\operatorname{H}^{s}(F\otimes_{R}M)_{\mathfrak{p}}$ , and that the above inequality is from Lemma 3.7.
@@
This implies a contradiction, and so $\mathfrak{p}=\mathfrak{m}$. As a result, Equation 3.3.3 becomes
@@
Let $s=n$. Since $\operatorname{H}^{0}(F)$ is finitely generated, every minimal generator in $\operatorname{H}^{0}(F)$ gives rise to a minimal generator in $\operatorname{H}^{0}(F)\otimes_{\operatorname{H}^{0}(R)}k$ by Nakayama’s Lemma. By [3, Lemma 3.2], $\operatorname{H}^{0}(F)\otimes_{\operatorname{H}^{0}(R)}k\cong\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)$ as $\sup(F)=0$. The minimal generator in $\operatorname{H}^{0}(F\otimes^{\mathbf{L}}_{R}k)$ gives rise to a nonzero element in $\operatorname{H}^{n}(F\otimes^{\mathbf{L}}_{R}(k\otimes_{R}M)$ by Diagram 3.3.2 in Theorem 3.5. The commutativity of Diagram 3.3.2 implies that there is a nonzero element in $\operatorname{H}^{n}(M\otimes_{R}F)$ that is annihilated by $I$, and so $\Gamma_{I}\operatorname{H}^{n}(M\otimes_{R}F)\neq 0$. Since $\inf(M\otimes_{R}F)=n$, this implies that $0pt_{R}(I,M\otimes_{R}F)=n$.
@@ 4.1. Definition
(4.1.1) $\operatorname{Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{n}((k\otimes^{\mathbf{L}}_{R}M)^{{\dagger}{\dagger}}),\operatorname{H}^{0}(E))\to\operatorname{Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{n}(M^{{\dagger}{\dagger}}),\operatorname{H}^{0}(E))$
which is nonzero since $\operatorname{H}^{0}(E)$ is a faithfully injective $\operatorname{H}^{0}(R)$-module. By [16, Thm. 4.10], Equation 4.1.1 can be rewritten as
(4.1.2) $\operatorname{H}^{-n}(\operatorname{\mathbf{R}Hom}_{R}((k\otimes_{R}M)^{{\dagger}{\dagger}},E))\to\operatorname{H}^{-n}(\operatorname{\mathbf{R}Hom}_{R}((M)^{{\dagger}{\dagger}},E))$
@@
The third isomorphism holds since $E\cong\Sigma^{d}\operatorname{\mathbf{R}\Gamma}_{\mathfrak{m}}(D)$ by [16, Prop. 7.25], and the fifth isomorphism holds by [16, Prop. 7.24] and Greenless-May duality. The sixth isomorphism holds since the DG-module $k$ is $\mathfrak{m}$-torsion. Equation 4.1.2 now becomes
(4.1.3) $\operatorname{H}^{-n}(\Sigma^{d}\operatorname{\mathbf{R}Hom}_{R}(k,M^{{\dagger}}))\to\operatorname{H}^{-n}(\Sigma^{d}(\operatorname{\mathbf{R}\Gamma}_{\mathfrak{m}}(M^{{\dagger}})))$
@@ 5. The Canonical Element Theorem for DG-Rings
(5.0.1) $\phi:\operatorname{\mathbf{R}Hom}_{\operatorname{H}^{0}(R)}(\operatorname{H}^{0}(R)\otimes^{\mathbf{L}}_{R}k,\operatorname{H}^{0}(D))\to\operatorname{\mathbf{R}Hom}_{\operatorname{H}^{0}(R)}(k,\operatorname{H}^{0}(D))$
@@
(5.0.2) $\xi_{\operatorname{H}^{0}(D)}^{d}:\operatorname{Ext}_{R}^{d}(k,\operatorname{H}^{0}(D))\to\operatorname{H}_{\mathfrak{m}}^{d}(\operatorname{H}^{0}(D))$
@@
Acknowledgments
The author wishes to thank his thesis advisor, Tom Marley, for both suggesting the topic behind this paper, and for the many discussions and suggestions that shaped this paper.